OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

@dan815 @Hero @ParthKohli @ganeshie8

OpenStudy (michele_laino):

hint: please you have to apply these formulas: \[\Large \begin{gathered} x = r\cos \theta \hfill \\ y = r\sin \theta \hfill \\ \\ {x^2} + {y^2} = {r^2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

using those formulas, you have to find the values for r and \theta

OpenStudy (anonymous):

so 2=rcostheta and -2=rsintheta